# ST-MambaSync: The Confluence of Mamba Structure and Spatio-Temporal Transformers for Precipitous Traffic Prediction

### v1

@@ Abstract
Balancing accuracy with computational efficiency is paramount in machine learning, particularly when dealing with high-dimensional data, such as spatial-temporal datasets. This study introduces ST-MambaSync, an innovative framework that integrates a streamlined attention layer with a simplified state-space layer. The model achieves competitive accuracy in spatial-temporal prediction tasks. We delve into the relationship between attention mechanisms and the Mamba component, revealing that Mamba functions akin to attention within a residual network structure. This comparative analysis underpins the efficiency of state-space models (also known as tensor models), elucidating their capability to deliver superior performance at reduced computational costs.
@@ 1 Introduction
Traffic prediction is essential for urban planning, road safety, and efficiency, significantly influencing the management of traffic flows, reduction of congestion, and the proactive deployment of resources to address potential disruptions. Accurate traffic forecasts allow for better traffic management and routing, which can reduce the risk of accidents and congestion. If commuters are informed about potential delays, they can take alternative routes or adjust their travel times, enhancing road safety and convenience. However, the challenge lies in achieving high accuracy but with low computational cost in both short-term (accurate predictions within a 1 to 5-minute timeframe) (Nagy and Simon, 2018) and especially in long-term (forecasting window to more than 30 minutes) prediction tasks.
Traditional approaches to traffic flow prediction, such as historical average (HA) (Smith, 1995), autoregressive integrated moving average (ARIMA) (Kumar and Vanajakshi, 2015), and support-vector regression (SVR) models (Wu et al., 2004), have laid the foundation for this field of study. While these models have provided valuable insights, their inherent simplicity and linear nature often limit their capacity to fully capture the nonlinear and high-dimensional relationships present in traffic data. As a result, the accuracy of these models may be compromised, particularly in complex and dynamic traffic scenarios.
@@
In recent years, the traffic flow prediction field has witnessed significant advancements with the introduction of deep learning techniques. While models such as Convolutional Neural Networks (CNNs) (Sayed et al., 2023) and Recurrent Neural Networks (RNNs) have shown promise in capturing spatial and temporal dependencies, they are not without limitations. CNNs excel in processing spatial information but may struggle with capturing long-range temporal dependencies. On the other hand, RNNs, including Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) networks, have demonstrated success in modeling temporal dynamics. However, RNNs can face challenges in capturing very long-term dependencies due to the vanishing gradient problem, which can hinder their ability to effectively model long sequences. Additionally, the sequential nature of RNNs can limit their parallelization capabilities, leading to slower training and inference times compared to feed-forward architectures like CNNs.
Recent research has explored the potential of Transformer models, originally developed for natural language processing tasks, in traffic flow prediction. Transformers introduce a novel approach to capturing dependencies through self-attention mechanisms. By allowing the model to weigh the importance of different parts of the input data, Transformers can effectively model complex relationships without the sequential processing limitations of RNNs. This ability to capture long-range dependencies has shown promise in improving the accuracy of long-term traffic flow predictions.
However, Transformer models come with their own set of challenges. The self-attention mechanism, while powerful, can be computationally expensive, particularly when dealing with large-scale traffic networks and long historical data sequences. The computational complexity of Transformers grows quadratically with the input sequence length, leading to slower training and testing phases compared to other models. Additionally, the high computational cost associated with Transformers can hinder their practical application in long-term traffic management systems, where swift predictions are crucial.
The Selective State of Space model (which is the so-called Mamba) (Gu and Dao, 2024) has a distinctive advantage in its capability to provide high-accuracy forecasts with reduced computational demands. This efficiency is particularly valuable in both short-term and long-term traffic management scenarios, where swift, reliable predictions are essential for effective congestion control, route optimization, and traffic flow regulation.
This paper introduces the Spatial-Temporal Selective State Space Model (ST-SSMs), an innovative framework for efficient and accurate traffic flow prediction. The ST-SSMs model seamlessly integrates spatial and temporal data through an ST-Mixer, eliminating the need to treat these data types separately. The integrated data is then processed by the ST-Mamba block, which excels at identifying and modeling the most significant patterns and dependencies in traffic data over extended periods.
A key feature of the ST-SSMs model is its ability to effectively and accurately process long-term traffic flow predictions by leveraging the strengths of Selective State Space Models (SSMs) within the ST-Mamba block. This makes ST-SSMs particularly effective for long-term analysis without compromising precision. As a state-of-the-art model designed specifically for non-graph-based Spatial-Temporal (ST) data, ST-SSMs introduce a new paradigm in traffic flow prediction. By utilizing the specially designed ST-Mamba block, the ST-SSMs model offers an advanced alternative to existing approaches, delivering more precise and faster forecasts of long-term traffic flow dynamics that reflect the intricate interplay of spatial and temporal factors.
@@
To the best of our knowledge, this paper presents the pioneering integration of a selective state-of-space model and attention that used in spatial temporal data and proves that mamba is a kind of attention with residual network. The key contributions of this research can be succinctly summarized as follows:
@@
  • This study introduces the integration of Mamba and attention blocks for handling spatial-temporal data in both real-time and long-term traffic forecasting tasks.
  • We provide theoretical proof that the Mamba can be conceptualized as an attention mechanism within a ResNet framework.
  • Through extensive experiments on real-world traffic datasets, we demonstrate the superior performance of our model compared to state-of-the-art benchmarks. Our approach achieves significant improvements in accuracy while reducing computation complexity.
2 Prelimanary
@@ 2.1 Attention
The self-attention mechanism, originally conceived for natural language processing (NLP), enriches the representation of feature data by revealing the underlying ’self-attention’ within the data set. Beginning with the input $X\in\mathbb{R}^{N\times d}$, the mechanism constructs queries ($Q$), keys ($K$), and values ($V$) via transformation matrices:
@@
Multi-head attention leverages multiple ’heads’ of $Q$, $K$, and $V$ to explore different representation subspaces, creating a rich, integrated output.
@@ 3 Method
Figure 1: Caption
@@ 3.1 Data Embedding
To encapsulate and reflect the dynamic temporal patterns present within the traffic data, we utilized a modifiable data embedding layer that processes the sequential input $\mathbf{X}_{t-M+1:t}$. Through the application of a dense neural layer, we extract the intrinsic feature embedding $\mathbf{Z}_{f}\in\mathbb{R}^{M\times N\times d_{e}}$:
$\mathbf{Z}_{f}=\text{Dense}(\mathbf{X}_{t-M+1:t})$ (10)
Here, $d_{e}$ signifies the dimension of the embedded features, and $\text{Dense}(\cdot)$ represents the applied dense layer. Furthermore, we introduce a parameterized dictionary for the embedding of weekdays $\mathbf{Z}_{w}\in\mathbb{R}^{7\times d_{e}}$ and another for the embedding of distinct times of the day $\mathbf{Z}_{h}\in\mathbb{R}^{288\times d_{e}}$, encapsulating the cyclical nature of weeks with 7 days and days with 288 time intervals. With $\mathbf{W}_{t}\in\mathbb{R}^{M}$ representing the weekday index and $\mathbf{H}_{t}\in\mathbb{R}^{M}$ representing the time-of-day index over the period from $t-M+1$ to $t$, we map these indices to their respective embeddings, yielding the weekday embedded features $\mathbf{Z}_{w_{t}}\in\mathbb{R}^{M\times d_{e}}$ and time-of-day embedded features $\mathbf{Z}_{h_{t}}\in\mathbb{R}^{M\times d_{e}}$. The combination and expansion of these embeddings generate the cyclical feature embedding $\mathbf{Z}_{c}\in\mathbb{R}^{M\times N\times 2d_{e}}$, which is utilized to incorporate periodic patterns into the traffic data.
Considering the rhythmic progression of time and the interlinked nature of traffic events, traffic sensors yield data with unique temporal traits. To address the need for a uniform approach to encapsulate these spatio-temporal dynamics, a shared spatio-temporal adaptive embedding, $\mathbf{Z}_{s}\in\mathbb{R}^{M\times N\times d_{s}}$, is put forth. This embedding is initialized utilizing Xavier uniform initialization, a technique that primes the model’s weights to avoid excessively large or small gradients initially, and thereafter, it is treated as a model parameter.
@@
$\mathbf{Z}=\textbf{Concatenate}(\mathbf{Z}_{f};\mathbf{Z}_{c};\mathbf{Z}_{s})$ (11)
@@ 3.3.1 ST-Mamba Layer
The ST-Mamba layer, described , leverages the discretization of continuous state-space models (SSM) as in Section 2.2. We represent the input to the ST-Mamba block as $\mathcal{H}$, where $\mathcal{H}=\text{LayerNorm}(\mathbf{\bar{X}})$. This representation is fed into the selective state-space model (SSM) layer for further processing. The linear transformation within the layer results in:
@@
which $\mathbf{U}\in\mathbb{R}^{d_{h}\times\mathcal{T}}$ denotes the hidden representation of the latent state at each iteration step $k$. The goal is to compute the output $\mathbf{Y}\in\mathbb{R}^{d_{inner}\times\mathcal{T}}$, and subsequently project $\mathbf{Y}$ back to the input shape $\mathcal{T}\times d_{h}$.
@@
The initialization of ST-SSM parameters is crucial for the functionality of the layer:
  • $\mathbf{A}\in\mathbb{R}^{d_{inner}\times d_{state}}$: Structured state transition matrix initialized using HiPPO to capture long-range dependencies.
  • $\mathbf{B}\in\mathbb{R}^{d_{state}\times\mathcal{T}}$: Computed as $\mathbf{B}=s_{B}(\mathbf{U})$, where $s_{B}(\cdot)$ is a learnable linear projection.
  • $\mathbf{C}\in\mathbb{R}^{d_{state}\times\mathcal{T}}$: Output projection matrix, computed as $\mathbf{C}=s_{C}(\mathbf{U})$, where $s_{C}(\cdot)$ is another learnable linear projection.
  • $\mathbf{D}\in\mathbb{R}^{d_{inner}}$: Learnable parameter that allows bypassing the state transformation for direct input-to-output information transfer.
  • $\Delta\in\mathbb{R}^{\mathbb{R}^{d_{state}\times\mathcal{T}}}$: Step size parameter calculated using $\Delta=\tau_{\Delta}(\text{Parameter}+s_{\Delta}(\mathbf{U}))$, with $\tau_{\Delta}$ as the softplus function and $s_{\Delta}(\cdot)$ as a linear projection.
@@
Discretizing the continuous-time parameters to obtain discrete-time SSM parameters involves:
@@
where $\exp(\cdot)$ denotes the matrix exponential, and $\mathbf{I}$ is the identity matrix of appropriate size. The discrete-time matrices, $\mathbf{\tilde{A}}$ and $\mathbf{\tilde{B}}$, are then used in the recurrence of the selective ST-Mamba layer:
@@
where $\odot$ represents the Hadamard product. The process iterates over each step from $k=1$ to $\mathcal{T}$, ensuring the transformation of each step’s data through the SSM layer. The final output $\mathbf{Y}$ is shaped into $\mathbb{R}^{\mathcal{T}\times d_{inner}}$ to match the input dimensions.

### v3

# ST-MambaSync: The Complement of Mamba and Transformers for Spatial-Temporal in Traffic Flow Prediction

@@ Abstract
Accurate traffic flow prediction is pivotal for optimizing traffic management, enhancing road safety, and reducing environmental impacts. Traditional models, however, struggle with the demands of processing long sequence data, requiring substantial memory and computational resources, and suffer from slow inference times due to the lack of a unified summary state. This paper introduces ST-MambaSync, a novel traffic flow prediction model that integrates transformer technology with the innovative ST-Mamba block, marking a significant advancement in the field. We are the first to implement the Mamba mechanism—essentially an attention mechanism combined with ResNet—within a transformer framework, significantly improving the model’s explainability and performance. ST-MambaSync not only addresses key challenges related to data length and computational efficiency but also sets new standards for accuracy and processing speed through extensive comparative analysis. This breakthrough has profound implications for urban planning and real-time traffic management, establishing a new benchmark in traffic flow prediction technology.
@@ 1 Introduction
Accurate traffic flow prediction is pivotal for optimizing traffic management, enhancing road safety, and reducing environmental impacts. Traditional models, however, struggle with the demands of processing long sequence data, requiring substantial memory and computational resources, and suffer from slow inference times due to the lack of a unified summary state. This paper introduces ST-MambaSync, a novel traffic flow prediction model that integrates transformer technology with the innovative ST-Mamba block, marking a significant advancement in the field. We are the first to implement the Mamba mechanism—essentially an attention mechanism combined with ResNet—within a transformer framework, significantly improving the model’s explainability and performance. ST-MambaSync not only addresses key challenges related to data length and computational efficiency but also sets new standards for accuracy and processing speed through extensive comparative analysis. This breakthrough has profound implications for urban planning and real-time traffic management, establishing a new benchmark in traffic flow prediction technology.
1.1 Traditional Approaches for Traffic Flow Predictions
The literature on traffic flow prediction has historically relied on several traditional methodologies that have been instrumental in laying the foundational understanding of this field. Early approaches include the Historical Average (HA) method (Smith, 1995), which predicts future traffic by averaging past data. While straightforward, this method often fails to capture sudden or atypical changes in traffic patterns.
Further developments led to the adoption of the Autoregressive Integrated Moving Average (ARIMA) model, introduced in traffic studies by (Kumar and Vanajakshi, 2015). ARIMA models enhance forecasting accuracy by considering past values and their errors; however, they inherently assume linear relationships among data points, which limits their effectiveness in handling the non-linear dynamics typical of traffic flows.
Another significant approach is the Support Vector Regression (SVR) model, utilized by (Wu et al., 2004) for traffic prediction. SVR has been preferred for its ability to handle non-linear data through the use of kernel functions. Despite its advantages, SVR, like its predecessors, struggles to manage the high-dimensional and complex relationships present in traffic datasets. In addition to these methods, the k-Nearest Neighbors (k-NN) algorithm has also been applied to traffic flow prediction (Zhao et al., 2020). The k-NN method predicts traffic by considering the k most similar historical traffic patterns. While it can capture non-linear relationships, it is sensitive to the choice of k and may struggle with high-dimensional data.
Moreover, researchers have explored the use of Kalman Filtering (KF) for traffic prediction (Guo et al., 2014). KF is a recursive algorithm that estimates the state of a system based on noisy measurements. It has been used to model the dynamic nature of traffic flows, but its performance can be affected by the quality of the initial estimates and the assumptions made about the system’s behavior.
While these traditional models have provided valuable insights and served as a stepping stone in the evolution of traffic flow prediction, their simplicity and linear assumptions often fall short in complex and dynamic traffic scenarios. This limitation is primarily due to their inability to fully capture the nonlinear interactions and dependencies inherent in traffic data, thus potentially compromising the accuracy of the predictions in more challenging environments. As traffic systems become increasingly complex, the need for more sophisticated, non-linear models that can effectively handle these complexities becomes evident, paving the way for the next generation of traffic prediction methodologies.
1.2 Deep Learning Method in Traffic Flow Prediction
@@
In recent years, the field of traffic flow prediction has seen notable progress with the integration of deep learning technologies. Deep learning models like Convolutional Neural Networks (CNNs) (Sayed et al., 2023) and Recurrent Neural Networks (RNNs) have been effective in understanding spatial and temporal patterns. CNNs are adept at handling spatial data but often falter with long-range temporal patterns. Conversely, RNNs, including variants like Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU), are good at temporal dynamics but struggle with very long-term dependencies due to issues like the vanishing gradient problem, which impedes their performance on lengthy sequences. Moreover, the inherently sequential processing of RNNs restricts their parallelization, slowing down both training and inference compared to CNNs.
Transformers, originally designed for natural language processing, have recently been adapted for traffic flow prediction, introducing a new method for dependency recognition via self-attention mechanisms. These mechanisms allow Transformers to evaluate the significance of different segments of input data, enabling them to handle complex dependencies more effectively than RNNs and without the need for sequential data processing. This capacity to manage long-range dependencies has shown potential in enhancing the accuracy of predictions for long-term traffic flows.
Despite their advantages, Transformers also face significant challenges. The self-attention mechanism, while effective, requires substantial computational resources, especially when applied to large traffic networks and extensive historical data. The computational demands of Transformers increase quadratically with the length of the input sequence, which can slow down both training and testing. The high computational needs may also limit the practicality of Transformers in real-time traffic management systems that require fast predictions.
1.3 State Space Model
Given the limitations of existing deep learning methods, the Selective State of Space model (commonly referred to as Mamba) (Gu and Dao, 2024) stands out for its ability to deliver high-accuracy forecasts while requiring less computational effort. This efficiency is particularly crucial in both short-term and long-term traffic management scenarios, where quick and reliable predictions are vital for effective congestion control, route optimization, and traffic regulation. A recent study (Shao et al., 2024) marked the first to apply the Mamba model to spatial-temporal traffic flow prediction, demonstrating promising results in reducing computational costs. However, there remains room for improvement in balancing accuracy with manageable computational demands.
This paper introduces the Spatial-Temporal Mamba Transformer (ST-MambaSync), a novel framework that efficiently integrates the popular transformer and Mamba methods for accurate traffic flow prediction. The ST-MambaSync model comprises two main components: the ST-Transformer and the ST-Mamba Block. The ST-Transformer efficiently processes data, capturing global information through spatial and temporal features. In contrast, the ST-Mamba Block, which includes an ST-Mixer, converts the tensor into a matrix. This matrix is then fed into the ST-Mamba layer, which updates individual hidden states and extends memory for long-range data, focusing more on local information. The combination on ST-Transformer and ST-Mamba Block not only enhances both global and local features but also accelerates computation, making it an effective component of our integrated approach for managing complex traffic data.
1.4 Contribution
@@
To the best of our knowledge, this paper introduces a groundbreaking integration of the selective state-of-space model (Mamba) with attention mechanisms specifically tailored for spatial-temporal data, establishing that Mamba effectively functions as a type of attention within a ResNet framework. The principal contributions of our research are summarized as follows:
@@
  • This study is the first to combine Mamba and attention blocks to manage spatial-temporal data, enhancing both real-time and long-term traffic forecasting.
  • We provide theoretical evidence showing that the Mamba model operates as an attention mechanism within a ResNet framework. This integration with the Transformer and the ST-Mamba block significantly boosts the model’s capacity to assimilate both comprehensive and granular information, thereby improving its performance in handling complex datasets.
  • Through rigorous testing on real-world traffic datasets, our model demonstrates superior performance relative to existing benchmarks, achieving notable gains in accuracy and efficiency while lowering computational demands.
This innovative approach not only advances the field of traffic flow prediction but also sets a new standard for the application of hybrid models in complex data environments.
2 Preliminary and Problem Statement
2.1 Notations
Road Network
In this study, we define a road network as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set $\mathcal{V}={v_{1},\ldots,v_{N}}$ corresponds to $N$ critical points such as traffic sensors and intersections on the roads. Each element of the edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents a section of road. Within this network, $\mathbf{X}_{t}\in\mathbb{R}^{N\times d}$ represents the traffic flow at the $N$ nodes at any given timestamp $t$, encapsulating flow characteristics across these nodes, with $d$ representing the dimensions of the features. Over a period $T>0$, the traffic data is aggregated into a mode-3 tensor $\mathbf{X}\in\mathbb{R}^{T\times N\times d}$, organizing the data along temporal, spatial, and feature-specific dimensions.
Problem Statement
The objective of traffic flow forecasting is to predict future traffic conditions accurately using historical data. For this purpose, we define a function $f$ within the context of a road network $\mathcal{G}$, which leverages traffic flow data from the past $M$ timestamps to forecast traffic conditions over the next $Z$ timestamps. This relationship can be mathematically formulated as:
$f\left([\mathbf{X}_{t-M+1},\ldots,\mathbf{X}_{t}];\mathcal{G}\right)\mapsto[\mathbf{X}_{t+1},\ldots,\mathbf{X}_{t+Z}],$
where $\mathbf{X}_{t}$ denotes the traffic flow tensor at time $t$. Throughout the model learning phase, we define $t$ to range from $M$ to $T-Z$ in order to efficiently utilize the observed traffic flow tensor $\mathbf{X}$.
@@ 2.2 Attention
The self-attention mechanism, originally conceived for natural language processing (NLP), enriches the representation of feature data by revealing the underlying “self-attention” within the data set. Beginning with the input $X\in\mathbb{R}^{N\times d}$, the mechanism constructs queries ($Q$), keys ($K$), and values ($V$) via transformation matrices:
@@
Multi-head attention leverages multiple “heads” of $Q$, $K$, and $V$ to explore different representation subspaces, creating a rich, integrated output.
@@ 3 Method
Figure 1: The framework of proposed ST-MambaSync.
@@ 3.1 Data Embedding
To encapsulate and reflect the dynamic temporal patterns present within the traffic data, we utilized a modifiable data embedding layer that processes the sequential input $\mathbf{X}_{t-M+1:t}$. Through the application of a dense neural layer, we extract the intrinsic feature embedding $\mathbf{Z}^{f}_{t}\in\mathbb{R}^{M\times N\times d_{e}}$:
$\mathbf{Z}^{f}_{t}=\text{Dense}(\mathbf{X}_{t-M+1:t})$ (12)
Here, $d_{e}$ signifies the dimension of the embedded features, and $\text{Dense}(\cdot)$ represents the applied dense layer. Furthermore, we introduce a parameterized dictionary for the embedding of weekdays $\mathbf{Z}_{w}\in\mathbb{R}^{7\times d_{e}}$ and another for the embedding of distinct times of the day $\mathbf{Z}_{h}\in\mathbb{R}^{288\times d_{e}}$, encapsulating the cyclical nature of weeks with 7 days and days with 288 time intervals. With $\mathbf{W}_{t}\in\mathbb{R}^{M}$ representing the weekday index and $\mathbf{H}_{t}\in\mathbb{R}^{M}$ representing the time-of-day index over the period from $t-M+1$ to $t$, we map these indices to their respective embeddings, yielding the weekday embedded features $\mathbf{Z}_{w_{t}}\in\mathbb{R}^{M\times d_{e}}$ and time-of-day embedded features $\mathbf{Z}_{h_{t}}\in\mathbb{R}^{M\times d_{e}}$. The combination and expansion of these embeddings generate the cyclical feature embedding $\mathbf{Z}^{c}_{t}\in\mathbb{R}^{M\times N\times 2d_{e}}$, which is utilized to incorporate periodic patterns into the traffic data.
Considering the rhythmic progression of time and the interlinked nature of traffic events, traffic sensors yield data with unique temporal traits. To address the need for a uniform approach to encapsulate these spatio-temporal dynamics, a shared spatio-temporal adaptive embedding, $\mathbf{Z}^{s}_{t}\in\mathbb{R}^{M\times N\times d_{s}}$, is put forth. This embedding is initialized utilizing Xavier uniform initialization, a technique that primes the model’s weights to avoid excessively large or small gradients initially, and thereafter, it is treated as a model parameter.
@@
$\mathbf{Z}_{t}=\textbf{Concatenate}(\mathbf{Z}^{f}_{t};\mathbf{Z}^{c}_{t};\mathbf{Z}^{s}_{t})$ (13)
@@ 3.3.1 ST-Mamba Layer
The ST-Mamba layer, as described in Section 2.3, utilizes the discretization of continuous state-space models (SSM). We denote the input to the ST-Mamba block as $\mathcal{H}$, which is obtained by applying LayerNorm to $\mathbf{\bar{X}}$. This processed input is then subjected to the selective state-space model (SSM) layer, where a linear transformation produces:
@@
with $\mathbf{U}\in\mathbb{R}^{d_{h}\times\mathcal{T}}$ representing the hidden state’s latent representation at each iteration step $k$. The objective is to calculate the output $\mathbf{Y}\in\mathbb{R}^{d_{inner}\times\mathcal{T}}$, which is then projected back to match the input’s dimensions $\mathcal{T}\times d_{h}$.
@@
Initializing parameters in the ST-SSM is vital for its operation:
  • $\mathbf{A}\in\mathbb{R}^{d_{inner}\times d_{state}}$: This structured state transition matrix is initialized using HiPPO to ensure capturing of long-range dependencies.
  • $\mathbf{B}\in\mathbb{R}^{d_{state}\times\mathcal{T}}$: Calculated as $\mathbf{B}=s_{B}(\mathbf{U})$, where $s_{B}(\cdot)$ is a learnable linear projection.
  • $\mathbf{C}\in\mathbb{R}^{d_{state}\times\mathcal{T}}$: Output projection matrix, derived as $\mathbf{C}=s_{C}(\mathbf{U})$, where $s_{C}(\cdot)$ is a learnable projection.
  • $\mathbf{D}\in\mathbb{R}^{d_{inner}}$: A learnable parameter that facilitates the direct transfer of information from input to output, bypassing the state transformation.
  • $\Delta\in\mathbb{R}^{d_{state}\times\mathcal{T}}$: Step size parameter, determined using $\Delta=\tau_{\Delta}(\text{Parameter}+s_{\Delta}(\mathbf{U}))$, with $\tau_{\Delta}$ being the softplus function and $s_{\Delta}(\cdot)$ a linear projection.
@@
The conversion of continuous-time parameters into discrete-time SSM parameters involves:
@@
where $\exp(\cdot)$ is the matrix exponential function, and $\mathbf{I}$ represents the identity matrix of suitable size. These discrete-time matrices, $\mathbf{\tilde{A}}$ and $\mathbf{\tilde{B}}$, facilitate the recurrence within the selective ST-Mamba layer:
@@
with $\odot$ denoting the Hadamard product. This iterative process spans each step from $k=1$ to $\mathcal{T}$, ensuring that each data step is transformed through the SSM layer. The final output, $\mathbf{Y}$, is reshaped into $\mathbb{R}^{\mathcal{T}\times d_{inner}}$ to align with the original input dimensions.